Crop & CanopyA life in growing
FIELD NOTES

Rice

Planting, care and harvest, with the changes along the way.

Read the growing guide ↗

  1. False smut in Vishnu Bhog

    Vishnu Bhog faced another setback, with spraying planned for the following day.

    Farm note · Observation
  2. Fallen rice changes the harvest decision

    Kala Namak and 1718 had fallen; we needed to choose how to bring them in.

    Farm note · Observation
  3. Jhanga removal in 1718 completed

    Manual removal of wild or volunteer paddy (Jhanga) from the 1718 fields was reported complete on 20 October.

    Farm note · Completed
  4. Navara Red Rice harvested

    Navara Red Rice has been harvested.

    Farm note · Completed
  5. Affected Sonam fields sprayed for leafroller

    The Sonam fields where Rice Leafroller had been identified have now been sprayed with insecticide.

    Farm note · Observation
  6. 1718 wild-paddy removal still ongoing

    Manual removal of wild or volunteer paddy, recorded as Jhanga, was still ongoing in 1718.

    Farm note · Ongoing
  7. Second leafroller spray round ongoing

    A second insecticide-spraying round for leafroller was ongoing.

    Farm note · Ongoing
  8. Significant Jhanga infestation; removal started

    A significant wild- or volunteer-paddy problem, recorded as Jhanga, was reported in 1718.

    Farm note · Ongoing
  9. Jhanga cleared; grain-filling spray ongoing

    Wild paddy/Jhanga removal in Vishnu Bhog was reported complete.

    Farm note · Ongoing
  10. Wild-paddy removal remains on the task list

    Removal of wild paddy/Jhanga remains an active crop task.

    Farm note · Ongoing
  11. Grain-filling spray remains planned

    The task list still called for a planned Vishnu Bhog grain-filling spray after Jhanga removal.

    Farm note · Planned
  12. Rogueing continues; grain-filling spray to follow

    Manual removal of Jhanga in rice area B remains ongoing.

    Farm note · Mixed progress
  13. Rice area B rogueing continues; area A removal deferred

    Manual removal of Jhanga (wild or volunteer paddy) is ongoing in rice area B.

    Farm note · Mixed progress
  14. Jhanga identified; manual removal starting

    Jhanga (wild or volunteer paddy from previous years) was identified in multiple fields, mainly in Vishnu Bhog.

    Farm note · Observation
  15. Jhanga also found in rice area A

    Jhanga was also identified in rice area A.

    Farm note · Planned
  16. Seasonal fertilisation completed across all paddy

    The final top-dressing and fertiliser applications were reported complete throughout the paddy crop.

    Farm note · Completed
  17. One paddy area finished; final dressing continued elsewhere

    The fertiliser programme was reported complete in one main rice area.

    Farm note · Partly complete
  18. Vishnu Bhog fertilisation complete

    The second top-dressing was completed at a reported 12 kg urea per bigha.

    Farm note · Completed
  19. 1718 and Sonam second rounds in progress

    Second top-dressing in the rice recorded as 1718 and Sonam was underway but delayed by rain.

    Farm note · Ongoing
  20. Leafroller treatment mostly completed

    Spraying was reported complete across the main paddy areas, including the direct-sown crop.

    Farm note · Partly complete
  21. Leafroller treatment continued

    Leafroller damage was reported beyond the first affected rice crop.

    Farm note · Ongoing
  22. Irrigation and leafroller scouting

    Irrigation had started in the main rice areas.

    Farm note · Mixed progress
  23. Direct-sown 1718 fertilised

    The separately direct-sown 1718 crop was reported fertilised after the wider August nutrient work.

    Farm note · Observation
  24. Nutrient work progressed by variety

    Vishnu Bhog and Navara were reported complete; Kala Namak was about 70% complete, while 1718 and Sonam were about half complete.

    Farm note · Partly complete
  25. Flood recovery and fertilisation

    A pond-side paddy area had drained and the crop survived.

    Farm note · Mixed progress
  26. Weed treatment and nutrients advanced

    Herbicide work was complete in one direct-sown section.

    Farm note · Mixed progress
  27. Floodwater pumped from a low area

    A waterlogged section needed pumping again; the removed water was reused for irrigation and transplant preparation in another part of the farm.

    Farm note · Observation
  28. Vishnu Bhog transplanting completed

    Transplanting of the Vishnu Bhog section that began in late July was reported complete.

    Farm note · Completed
  29. Heavy rain disrupted transplanting

    Flooding in a low section required drainage, while transplanting elsewhere continued as water and labour allowed.

    Farm note · Observation
  30. Vishnu Bhog transplanting began

    Vishnu Bhog transplanting began.

    Farm note · Ongoing
  31. 1718 transplanting began

    Transplanting began for 1718.

    Farm note · Ongoing
  32. Navara trial transplanted

    The small Navara red-rice trial was transplanted.

    Farm note · Observation
  33. 1718 direct broadcasting completed

    A separate block recorded as 1718 was broadcast directly rather than transplanted.

    Farm note · Completed
  34. Additional nurseries sown

    Nursery sowing included Sonam, Kala Namak, Navara red rice and another 1718 batch.

    Farm note · Observation
  35. Vishnu Bhog nursery broadcast

    Cleaned Vishnu Bhog seed was broadcast in the nursery in the late afternoon.

    Farm note · Observation
  36. Vishnu Bhog seed cleaned

    Seed winnowing was completed before nursery sowing; rejected material was separated from the clean seed.

    Farm note · Observation